# A Comprehensive Evaluation of Parameter-Efficient Fine-Tuning on Code Smell Detection

### v2

@@ 3.4. Base Model Selection
In this work, we aim to explore the effectiveness of PEFT methods for the task of code smell detection. To ensure a thorough analysis, we included both small and large LMs as our base models. Consistent with the definition adopted by Weyssow et al. (weyssow2023exploring, ), we classified models with 1B or more parameters as LLMs, and those with fewer parameters as SLMs.
Table 4. Selected LMs and their corresponding parameter size and architecture
Model #Params Architecture
SLMs CodeBERT 125M Encoder-only
GraphCodeBERT 125M
UnixCoder 126M
CodeT5 220M Encoder-decoder
LLMs StarCoderBase-1B 1.1B Decoder-only
StarCoderBase-3B 3B
DeepSeek-Coder-1.3B 1.3B
DeepSeek-Coder-6.7B 6.7B
CodeLlama-7B 7B
@@
Based on the above criteria, we employed 4 SLMs (i.e., CodeBERT (feng2020codebert, ), GraphCodeBERT (guo2021graphcodebert, ), UnixCoder (guo2022unixcoder, ), and CodeT5 (wang2021codet5, )) and 5 LLMs (StarCoderBase-1B (li2023starcoder, ), StarCoderBase-3B (li2023starcoder, ), DeepSeek-Coder-1.3B (guo2024deepseekcoder, ), DeepSeek-Coder-6.7B (guo2024deepseekcoder, ), and CodeLlama-7B (roziere2024codellama, )) for our experiments. To ensure a comprehensive analysis, we deliberately included models representing all three major Transformer architectures: encoder-only, encoder-decoder, and decoder-only. Although decoder-only models were originally designed for generative tasks (vaswani2017attention, ; radford2019language, ), prior studies have shown their competitive performance in sequence classification tasks as well (wolf2020transformers, ; kariyappa2024progressive, ). Therefore, we also included decoder-only models as base models in our study. Additionally, all the models we selected have demonstrated encouraging results across various code understanding tasks like code clone detection, defect detection, etc. (liu2024vul, ; karmakar2021what, ; lu2021codexglue, ; li2022automating, ; tang2023csgvd, ; zeng2022an, ; chen2023diversevul, ; wang2023one, ; fu2022vulrepair, ; nguyen2024empirical, ; su2024distilled, ). Table 4 presents the details of the 9 base models, including their parameter sizes and architectural types.
@@ 3.5. Metrics
We use metrics that align with previous research (sharma2021code, ; ho2023fusion, ; pecorelli2019comparing, ) to maintain consistency in evaluating the effectiveness of PEFT methods for code smell detection on LLMs. Specifically, we utilize precision, recall, F1-score, and MCC (Matthews Correlation Coefficient) for assessing the performance of detecting CC and CM. We include MCC because, unlike other metrics like F1-score, MCC accounts for true negative instances (yao2020assessing, ). These metrics are calculated using the following formulas:
@@
[…]
@@
where $TP$, $TN$, $FP$, $FN$ represent true positives, true negatives, false positives, and false negatives, respectively.
@@ 3.6. Implementation Details
The code for our experiments is implemented by the PEFT library [^4] https://huggingface.co/docs/peft/index from Hugging Face. After reviewing prior studies (liu2023empirical, ; liu2024delving, ; li2024comprehensive, ; wang2023one, ; weyssow2023exploring, ) and observing the results of pilot experiments, we set the learning rate to $3e^{-4}$ for all PEFT methods and $1e^{-5}$ for full fine-tuning the SLMs. Throughout all the fine-tuning experiments, we utilized the AdamW optimizer (loshchilov2017decoupled, ) and fixed the Transformer input window size at 1,024. The choice to limit the input length to 1,024 tokens was driven by hardware constraints. As the GPU memory consumption of transformer-based models grows roughly quadratically with input length (tay2022efficient, ), this setting enables all selected base models to train within available GPU memory capacity. All the models were trained for 10 epochs, and we selected the epochs at which models exhibit minimum loss on the validation sets for inference.
Table 5 shows the hyperparameter settings of different PEFT methods. We adopted a batch size of 1 for all experiments to avoid GPU out-of-memory issues. We used a fixed random seed (42) across all runs to ensure experimental reproducibility. For prompt tuning and prefix tuning, we used 20 virtual tokens, which are learnable vectors prepended to the input and optimized during training. This setting is widely used in the literature and offers a balance between representational capacity and computational cost (li2021prefix, ; weyssow2023exploring, ). As for LoRA, we set the rank $r$ to 8, the scaling factor $\alpha$ to 16, and the dropout rate to 0.1 to prevent overfitting, following configurations reported in (liu2024delving, ). For both LoRA and ${IA}^{3}$, we applied parameter-efficient adaptation to the core components of the Transformer architecture, including the query ($Q$), key ($K$), and value ($V$) projections in the attention layers. In the case of $(\text{IA})^{3}$, we also included the Feed-Forward Network (FFN) for adaptation (see Section 2.3). We executed all models on a single Linux GPU server equipped with a 40GB NVIDIA A100 GPU. Additionally, we have released our datasets and source code online for replication purpose (replicationPackage, ) where more details can be found.
Table 5. Hyperparameter settings of different PEFT methods and full fine-tuning
PEFT Method Hyperparameter Setting Batch Size Seed
@@
Full Fine-tuing None
@@ 4. Results
In this section, we first illustrate the experiment setup for each RQ formulated in Section 3, and we then present the results and findings for the four RQs.
@@ 4.1.1. Experiment Setup
SLMs CodeBERT Full FT 70.54% 44.41% 80.57% 61.19%
Prompt Tuning 62.38% 25.90% 76.13% 54.00%
Prefix Tuning 63.88% 27.87% 81.33% 62.94%
LoRA 73.04% 46.10% 71.00% 49.79%
$(IA)^{3}$ 69.44% 40.49% 81.32% 62.96%
GraphCodeBERT Full FT 74.01% 48.08% 80.66% 61.33%
Prompt Tuning 65.24% 30.81% 78.90% 58.60%
Prefix Tuning 69.82% 39.97% 79.18% 60.46%
LoRA 73.15% 46.30% 80.03% 60.93%
$(IA)^{3}$ 72.50% 45.00% 80.00% 60.45%
UnixCoder Full FT 73.04% 46.10% 82.39% 64.89%
Prompt Tuning 69.35% 38.70% 78.66% 57.95%
Prefix Tuning 71.63% 43.26% 82.71% 65.84%
LoRA 72.19% 46.04% 73.18% 51.52%
$(IA)^{3}$ 73.47% 46.98% 82.23% 64.47%
CodeT5 Full FT 73.01% 46.19% 81.13% 62.74%
LoRA 33.33% 0.00% 83.29% -2.95%
$(IA)^{3}$ 43.38% -4.05% 40.32% 16.70%
LLMs StarCoderBase-1B Prompt Tuning 71.07% 42.21% 80.05% 60.11%
Prefix Tuning 70.10% 40.23% 76.88% 54.01%
LoRA 41.41% -1.62% 45.41% 7.01%
$(IA)^{3}$ 74.27% 49.01% 80.90% 62.01%
StarCoderBase-3B Prompt Tuning 69.12% 38.29% 81.87% 63.85%
Prefix Tuning 67.81% 35.69% 78.25% 56.92%
LoRA 40.13% 7.93% 68.41% 46.00%
$(IA)^{3}$ 74.68% 49.96% 82.07% 64.81%
Deepseek-Coder-1.3B Prompt Tuning 70.21% 40.44% 81.77% 63.75%
Prefix Tuning 69.38% 39.62% 79.06% 58.36%
LoRA 69.44% 40.49% 78.42% 57.96%
$(IA)^{3}$ 72.50% 45.01% 81.87% 63.82%
Deepseek-Coder-6.7B Prompt Tuning 67.93% 35.87% 81.19% 63.11%
Prefix Tuning 68.80% 37.62% 80.09% 60.57%
LoRA 69.98% 40.05% 81.34% 63.63%
$(IA)^{3}$ 70.84% 42.36% 83.34% 66.88%
CodeLlama-7B Prompt Tuning 68.68% 37.92% 81.25% 63.43%
Prefix Tuning 53.41% 15.95% 80.77% 62.08%
LoRA 70.12% 40.70% 80.70% 62.59%
$(IA)^{3}$ 70.48% 42.32% 83.00% 66.17%
@@
SLMs CodeBERT Full FT 74.09% 48.80% 81.98% 65.60%
Prompt Tuning 56.17% 13.07% 81.56% 64.57%
Prefix Tuning 62.82% 25.93% 81.81% 65.43%
LoRA 67.44% 36.17% 83.17% 68.48%
$(IA)^{3}$ 66.49% 38.48% 83.30% 68.18%
GraphCodeBERT Full FT 77.40% 54.95% 82.66% 67.42%
Prompt Tuning 66.88% 37.49% 83.75% 68.95%
Prefix Tuning 74.19% 48.39% 89.26% 79.26%
LoRA 72.23% 46.36% 85.85% 72.40%
$(IA)^{3}$ 69.84% 45.48% 85.58% 72.56%
UnixCoder Full FT 77.04% 56.76% 83.51% 69.49%
Prompt Tuning 77.32% 55.30% 83.68% 68.74%
Prefix Tuning 75.80% 51.64% 83.89% 69.28%
LoRA 70.97% 41.94% 83.60% 68.65%
$(IA)^{3}$ 73.95% 49.32% 83.96% 69.44%
CodeT5 Full FT 66.49% 38.48% 83.46% 69.12%
LoRA 54.07% 10.02% 70.03% 49.06%
$(IA)^{3}$ 54.65% 9.76% 46.79% -2.66%
LLMs StarCoderBase-1B Prompt Tuning 72.23% 46.36% 84.66% 71.21%
Prefix Tuning 75.65% 52.30% 85.44% 72.19%
LoRA 78.58% 60.68% 79.30% 58.85%
$(IA)^{3}$ 73.95% 49.32% 84.96% 71.71%
StarCoderBase-3B Prompt Tuning 68.96% 39.74% 83.63% 68.47%
Prefix Tuning 67.61% 35.78% 86.56% 74.06%
LoRA 70.85% 42.29% 83.70% 68.64%
$(IA)^{3}$ 69.35% 38.73% 84.90% 71.46%
DeepSeek-Coder-1.3B Prompt Tuning 78.98% 58.34% 84.17% 69.99%
Prefix Tuning 66.88% 37.49% 82.44% 64.90%
LoRA 78.98% 58.34% 82.72% 67.68%
$(IA)^{3}$ 83.87% 67.74% 85.84% 72.57%
DeepSeek-Coder-6.7B Prompt Tuning 78.89% 58.83% 85.56% 72.72%
Prefix Tuning 80.56% 61.81% 83.03% 68.14%
LoRA 77.32% 55.30% 83.43% 68.64%
$(IA)^{3}$ 77.32% 55.30% 85.34% 72.34%
CodeLlama-7B Prompt Tuning 72.52% 45.37% 85.49% 72.55%
Prefix Tuning 63.57% 30.67% 82.30% 64.61%
LoRA 77.32% 55.30% 83.10% 68.31%
$(IA)^{3}$ 74.19% 48.39% 83.54% 68.45%

### v3

@@ 3.6. Data Split
4. Experimental Setup
4.1. Task Formulation
We formulate code smell detection as a binary classification task. Given a code snippet $\mathcal{C}={c_{1},c_{2},\dots,c_{n}}$ consisting of a sequence of tokens, the goal is to learn a mapping function $f:\mathcal{C}\to\mathcal{Y}$, where $\mathcal{Y}\in{0,1}$. Here, $1$ indicates the presence of a specific code smell, and 0 denotes its absence.
As illustrated in Figure 5, we employ a Language Model (LM) as the backbone for this function $f$. To adapt the LM for this specific classification task, we append a sequence classification head on top of the model and optimize it using four Parameter-Efficient Fine-Tuning (PEFT) techniques.
4.2. Research Questions
RQ1: What is the effectiveness of different PEFT methods and full fine-tuning on various SLMs and LLMs? Previous studies have shown remarkable results using PEFT methods to fine-tune SLMs for software engineering tasks like code clone detection and defect detection (liu2023empirical; liu2024delving; li2024comprehensive; wang2023one; weyssow2023exploring). However, these studies have not paid attention to the effectiveness of PEFT methods for code smell detection. With this RQ, we aim to compare and validate the effectiveness of PEFT methods across multiple LMs of varying parameter sizes (see Section 4.3) specifically for code smell detection.
RQ2: What is the effectiveness of PEFT methods with different hyper-parameters? When fine-tuning LMs for downstream tasks, the configuration of hyper-parameters plays a crucial role in determining the final results by affecting the number of trainable parameters. In prompt tuning and prefix tuning, the number of virtual tokens serves as a critical hyper-parameter (lester2021power; li2021prefix). For LoRA, the key hyper-parameter is the rank $r$ (hu2022lora), while in $(IA)^{3}$, it is the selection of activated modules (liu2022few). Due to our limited computational resources and the popularity of LoRA compared to other PEFT methods (gupta2024beyond; zhang2024star), we decided to employ LoRA to fine-tune a SLM and an LLM with different $r$ values in a range of {8, 32, 64, 640, 1,280, 2,560} to explore the effect of hyper-parameter settings on PEFT methods for code smell detection.
RQ3: What is the effectiveness of PEFT methods in low-resource scenarios? Low-resource scenarios, characterized by constraints on the amount of training data, pose significant challenges to fine-tuning LMs for specific tasks. Due to the scarcity of real-world datasets (zakeri2023systematic), these scenarios are common in tasks such as code smell detection. Therefore, it is crucial to investigate how the number of training samples influences the effectiveness of PEFT methods for code smell detection.
RQ4: What is the effectiveness of PEFT methods compared to state-of-the-art code smell detection approaches? Existing code smell detectors include traditional heuristics-based and DL-based methods. Additionally, the use of general-purpose LLMs with In-Context Learning (ICL) has become increasingly popular in SE tasks (zhou2024large; shin2023prompt; li2025large; xu2024unilog), and such models can also be leveraged for code smell detection. With this RQ, we aim to assess how LMs utilizing PEFT methods compare against heuristics-based code smell detectors, established DL techniques, and general-purpose LLMs under zero- and few-shot settings for the task of code smell detection.
@@ 4.3. Base Model Selection
In this work, we aim to explore the effectiveness of PEFT methods for the task of code smell detection. To ensure a thorough analysis, we included both small and large LMs as our base models. Consistent with the definition adopted by Weyssow et al. (weyssow2023exploring), we classified models with 1B or more parameters as LLMs, and those with fewer parameters as SLMs.
Table 4. Selected LMs and their corresponding parameter sizes, context window limits, and architecture types
Model #Params Context Window Architecture
SLMs CodeBERT 125M 512 Encoder-only
GraphCodeBERT 125M 512
UnixCoder 125M 1,024
CodeT5 220M 512 Encoder-decoder
LLMs StarCoderBase-1B 1.1B 1,024 Decoder-only
StarCoderBase-3B 3B 1,024
DeepSeek-Coder-1.3B 1.3B 1,024
DeepSeek-Coder-6.7B 6.7B 1,024
CodeLlama-7B 6.7B 1,024
@@
Based on the above criteria, we employed 4 SLMs (i.e., CodeBERT (feng2020codebert), GraphCodeBERT (guo2021graphcodebert), UnixCoder (guo2022unixcoder), and CodeT5 (wang2021codet5)) and 5 LLMs (StarCoderBase-1B (li2023starcoder), StarCoderBase-3B (li2023starcoder), DeepSeek-Coder-1.3B (guo2024deepseekcoder), DeepSeek-Coder-6.7B (guo2024deepseekcoder), and CodeLlama-7B (roziere2024codellama)) for our experiments. To ensure a comprehensive analysis, we deliberately included models representing all three major Transformer architectures: encoder-only, encoder-decoder, and decoder-only. Although decoder-only models were originally designed for generative tasks (vaswani2017attention; radford2019language), prior studies have shown their competitive performance in sequence classification tasks as well (wolf2020transformers; kariyappa2024progressive). Therefore, we also included decoder-only models as base models in our study. Additionally, all the models we selected have demonstrated encouraging results across various code understanding tasks like code clone detection, defect detection, etc. (liu2024vul; karmakar2021what; lu2021codexglue; li2022automating; tang2023csgvd; zeng2022an; chen2023diversevul; wang2023one; fu2022vulrepair; nguyen2024empirical; su2024distilled). Table 4 presents the details of the 9 base models, including their parameter sizes, context window limits, and architectural types. The input sequences of CodeBERT, GraphCodeBERT, and CodeT5 were truncated at 512 tokens, which corresponds to their maximum supported input length. For all other models, inputs were uniformly truncated at 1,024 tokens.
4.4. Baselines
In RQ4, to comprehensively evaluate the effectiveness of PEFT-tuned LMs for code smell detection, we compare them against three distinct categories of baselines: traditional heuristics-based detectors, existing Deep Learning (DL)-based approaches, and state-of-the-art general-purpose LLMs using In-Context Learning (ICL).
4.4.1. Heuristics-based detectors
We select two prominent heuristics-based detectors that rely on predefined metrics and rule-based algorithms to identify code smells:
  • DesigniteJava (designitejava): An industrial-strength tool that detects architecture, design, and implementation smells by analyzing source code structure and calculating metrics (e.g., cyclomatic complexity and class coupling). In this study, we deploy DesigniteJava to detect CC, CM, and FE.
  • PMD (pmd): A widely adopted static analyzer that flags potential inefficiencies and bad practices in Java code through rigorous rule matching. In this study, we specifically task PMD with identifying DC instances.
These tools serve as foundational reference points in our study and are extensively cited in prior literature (sharma2021qscored; shah2023mining; barbosa2020revealing; fernandes2016review; aniche2018code; wu2024ismell). Furthermore, since these tools were utilized in our initial code smell collection phase (see Section 3.3), comparing against them provides critical insight into how much improvement our PEFT-based approach offers over the original labeling sources.
4.4.2. DL-based Approaches
We incorporate two representative DL models as baselines to evaluate the performance of established neural architectures:
  • DeepSmells (ho2023fusion): A model that utilizes a 1D-Convolutional Neural Network (1D-CNN) followed by a Long Short-Term Memory (LSTM) network to learn representations from tokenized source code sequences.
  • AE-Dense (sharma2021code): An Autoencoder-based model designed to learn dense vector representations of code instances in an unsupervised manner.
Collectively, these comparative DL models represent the state-of-the-art in leveraging sequential patterns and latent feature learning for code smell detection, respectively.
4.4.3. LLMs with In-Context Learning (ICL)
To benchmark the performance of our fine-tuned models against the “prompting” paradigm without parameter updates for code smell detection, we leverage two powerful general-purpose LLMs. These models exemplify the latest advances in commercial and open-weight architectures under the ICL paradigm:
  • GPT-4o-mini (hurst2024gpt4o): One of the most advanced proprietary LLMs developed by OpenAI, known for its superior reasoning capabilities and code understanding in lightweight deployments.
  • DeepSeek-v3 (liu2024deepseek): A current cutting-edge open-source mixture-of-experts (MoE) model (671B parameters), which serves as a strong reference for the performance of large-scale open-weight models.
Both models have demonstrated exceptional capabilities across natural language and programming tasks (rasheed2024task; bruni2025benchmarking; liu2024deepseek; jahin2025unveiling). We assess their effectiveness in code smell detection under three ICL settings: zero-shot, few-shot, and a criteria-informed setting that provides explicit labeling rules (refer to Section 5.4 for details).
@@ 4.5. Metrics
We use metrics that align with previous research (sharma2021code; ho2023fusion; pecorelli2019comparing) to maintain consistency in evaluating the effectiveness of PEFT methods for code smell detection on LLMs. Specifically, we utilize precision, recall, F1-score, and MCC (Matthews Correlation Coefficient) for assessing the performance of detecting CC and CM. We include MCC because, unlike other metrics like F1-score, MCC accounts for true negative instances (yao2020assessing). These metrics are calculated using the following formulas:
@@
(5) $\text{F1-Score}=2\times\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$
@@
where $TP$, TNTN, FPFP, FNFN represent true positives, true negatives, false positives, and false negatives, respectively.
@@ 4.6. Implementation Details
The code for our experiments is implemented by the PEFT library [^4] https://huggingface.co/docs/peft/index from Hugging Face. After reviewing prior studies (liu2023empirical; liu2024delving; li2024comprehensive; wang2023one; weyssow2023exploring) and observing the results of pilot experiments, we set the learning rate to $3e^{-4}$ for all PEFT methods and $1e^{-5}$ for full fine-tuning the SLMs. Throughout all the fine-tuning experiments, we utilized the AdamW optimizer (loshchilov2017decoupled) and fixed the Transformer input window size at 1,024. The choice to limit the input length to 1,024 tokens was driven by hardware constraints. As the GPU memory consumption of transformer-based models grows roughly quadratically with input length (tay2022efficient), this setting enables all selected base models to train within available GPU memory capacity. All the models were trained for 10 epochs, and we selected the epochs at which models exhibit minimum loss on the validation sets for inference.
Table 5 shows the hyper-parameter settings of different PEFT methods. Due to resource constraints, the batch size was set to 1 for LLMs with large parameter sizes to avoid out-of-memory issues. To ensure consistency and comparability, we adopted a batch size of 1 for all experiments. We used a fixed random seed (42) across all runs to ensure experimental reproducibility. For prompt tuning and prefix tuning, we used 20 virtual tokens, which are learnable vectors prepended to the input and optimized during training. This setting is widely used in the literature and offers a balance between representational capacity and computational cost (li2021prefix; weyssow2023exploring). As for LoRA, we set the rank $r$ to 8, the scaling factor $\alpha$ to 16, and the dropout rate to 0.1 to prevent overfitting, following configurations reported in (liu2024delving). For both LoRA and ${IA}^{3}$, we applied parameter-efficient adaptation to the core components of the Transformer architecture, including the query ($Q$), key ($K$), and value ($V$) projections in the attention layers. In the case of $(\text{IA})^{3}$, we also included the Feed-Forward Network (FFN) for adaptation (see Section 2.3.2). We executed all models on a single Linux GPU server equipped with a 40GB NVIDIA A100 GPU. Additionally, we have released our datasets and source code online for replication purpose (replicationPackage) where more details can be found.
Table 5. Hyper-parameter settings of different PEFT methods and full fine-tuning
PEFT Method Hyper-parameter Setting Batch Size Seed
@@
Full Fine-tuning None
@@ 5. Results
In this section, we present the results and findings for the four RQs formulated in Section 4.2. For each RQ, we first specify the particular experimental configuration and execution strategy based on the general setup in Section 4, followed by a detailed analysis of the results.
@@
SLMs CodeBERT Full FT 125M 70.54% 44.41% 80.57% 61.19%
Prompt Tuning 0.61M 62.38% 25.90% 76.13% 54.00%
Prefix Tuning 0.96M 63.88% 27.87% 81.33% 62.94%
LoRA 1.03M 73.04% 46.10% 71.00% 49.79%
$(IA)^{3}$ 0.67M 69.44% 40.49% 81.32% 62.96%
GraphCodeBERT Full FT 125M 74.01% 48.08% 80.66% 61.33%
Prompt Tuning 0.61M 65.24% 30.81% 78.90% 58.60%
Prefix Tuning 0.96M 69.82% 39.97% 79.18% 60.46%
LoRA 1.03M 73.15% 46.30% 80.03% 60.93%
$(IA)^{3}$ 0.67M 72.50% 45.00% 80.00% 60.45%
UnixCoder Full FT 125M 73.04% 46.10% 82.39% 64.89%
Prompt Tuning 0.61M 69.35% 38.70% 78.66% 57.95%
Prefix Tuning 0.96M 71.63% 43.26% 82.71% 65.84%
LoRA 1.03M 72.19% 46.04% 73.18% 51.52%
$(IA)^{3}$ 0.67M 73.47% 46.98% 82.23% 64.47%
CodeT5 Full FT 220M 73.01% 46.19% 81.13% 62.74%
LoRA 1.33M 33.33% 0.00% 83.29% -2.95%
$(IA)^{3}$ 0.16M 43.38% -4.05% 40.32% 16.70%
LLMs StarCoderBase-1B Prompt Tuning 0.05M 71.07% 42.21% 80.05% 60.11%
Prefix Tuning 1.97M 70.10% 40.23% 76.88% 54.01%
LoRA 3.59M 41.41% -1.62% 45.41% 7.01%
$(IA)^{3}$ 0.31M 74.27% 49.01% 80.90% 62.01%
StarCoderBase-3B Prompt Tuning 0.06M 69.12% 38.29% 81.87% 63.85%
Prefix Tuning 4.06M 67.81% 35.69% 78.25% 56.92%
LoRA 7.38M 40.13% 7.93% 68.41% 46.00%
$(IA)^{3}$ 0.62M 74.68% 49.96% 82.07% 64.81%
Deepseek-Coder-1.3B Prompt Tuning 0.05M 70.21% 40.44% 81.77% 63.75%
Prefix Tuning 1.97M 69.38% 39.62% 79.06% 58.36%
LoRA 2.36M 69.44% 40.49% 78.42% 57.96%
$(IA)^{3}$ 0.28M 72.50% 45.01% 81.87% 63.82%
Deepseek-Coder-6.7B Prompt Tuning 0.09M 67.93% 35.87% 81.19% 63.11%
Prefix Tuning 5.25M 68.80% 37.62% 80.09% 60.57%
LoRA 6.30M 69.98% 40.05% 81.34% 63.63%
$(IA)^{3}$ 0.75M 70.84% 42.36% 83.34% 66.88%
CodeLlama-7B Prompt Tuning 0.09M 68.68% 37.92% 81.25% 63.43%
Prefix Tuning 5.25M 53.41% 15.95% 80.77% 62.08%
LoRA 6.30M 70.12% 40.70% 80.70% 62.59%
$(IA)^{3}$ 0.75M 70.48% 42.32% 83.00% 66.17%
@@
SLMs CodeBERT Full FT 125M 74.09% 48.80% 81.98% 65.60%
Prompt Tuning 0.61M 56.17% 13.07% 81.56% 64.57%
Prefix Tuning 0.96M 62.82% 25.93% 81.81% 65.43%
LoRA 1.03M 67.44% 36.17% 83.17% 68.48%
$(IA)^{3}$ 0.67M 66.49% 38.48% 83.30% 68.18%
GraphCodeBERT Full FT 125M 77.40% 54.95% 82.66% 67.42%
Prompt Tuning 0.61M 66.88% 37.49% 83.75% 68.95%
Prefix Tuning 0.96M 74.19% 48.39% 89.26% 79.26%
LoRA 1.03M 72.23% 46.36% 85.85% 72.40%
$(IA)^{3}$ 0.67M 69.84% 45.48% 85.58% 72.56%
UnixCoder Full FT 125M 77.04% 56.76% 83.51% 69.49%
Prompt Tuning 0.61M 77.32% 55.30% 83.68% 68.74%
Prefix Tuning 0.96M 75.80% 51.64% 83.89% 69.28%
LoRA 1.03M 70.97% 41.94% 83.60% 68.65%
$(IA)^{3}$ 0.67M 73.95% 49.32% 83.96% 69.44%
CodeT5 Full FT 220M 66.49% 38.48% 83.46% 69.12%
LoRA 1.33M 54.07% 10.02% 70.03% 49.06%
$(IA)^{3}$ 0.16M 54.65% 9.76% 46.79% -2.66%
LLMs StarCoderBase-1B Prompt Tuning 0.05M 72.23% 46.36% 84.66% 71.21%
Prefix Tuning 1.97M 75.65% 52.30% 85.44% 72.19%
LoRA 3.59M 78.58% 60.68% 79.30% 58.85%
$(IA)^{3}$ 0.31M 73.95% 49.32% 84.96% 71.71%
StarCoderBase-3B Prompt Tuning 0.06M 68.96% 39.74% 83.63% 68.47%
Prefix Tuning 4.06M 67.61% 35.78% 86.56% 74.06%
LoRA 7.38M 70.85% 42.29% 83.70% 68.64%
$(IA)^{3}$ 0.62M 69.35% 38.73% 84.90% 71.46%
DeepSeek-Coder-1.3B Prompt Tuning 0.05M 78.98% 58.34% 84.17% 69.99%
Prefix Tuning 1.97M 66.88% 37.49% 82.44% 64.90%
LoRA 2.36M 78.98% 58.34% 82.72% 67.68%
$(IA)^{3}$ 0.28M 83.87% 67.74% 85.84% 72.57%
DeepSeek-Coder-6.7B Prompt Tuning 0.09M 78.89% 58.83% 85.56% 72.72%
Prefix Tuning 5.25M 80.56% 61.81% 83.03% 68.14%
LoRA 6.30M 77.32% 55.30% 83.43% 68.64%
$(IA)^{3}$ 0.75M 77.32% 55.30% 85.34% 72.34%
CodeLlama-7B Prompt Tuning 0.09M 72.52% 45.37% 85.49% 72.55%
Prefix Tuning 5.25M 63.57% 30.67% 82.30% 64.61%
LoRA 6.30M 77.32% 55.30% 83.10% 68.31%
$(IA)^{3}$ 0.75M 74.19% 48.39% 83.54% 68.45%